OpenStudy (anonymous):

Chemistry help?!?! A calorimeter contains 100 g of water at 25°C. You heat a 100 g copper block to its melting point (1,358 K) and plunge it in the calorimeter. Assume that the final temperature of the block reaches 25°C. How much heat was released from the copper block? At the end of the experiment, how many grams of water are liquid and at what temperature? How many grams of water are steam and at what temperature? Show your work. Specific heat of liquid water = 4.18 J/g–°C Specific heat of copper = 0.385 J/g–°C ΔHv for water = 40.7 kJ/mol

OpenStudy (aaronq):

have you tried anything?

OpenStudy (anonymous):

No i am lost, do you know anything that could help me get the answers, could you do a step by step explanation with the answer?

OpenStudy (aaronq):

you can start by determining the heat, q, released from the copper block. q=mCdt m= mass C=specific heat capacity dT= change in temp

OpenStudy (anonymous):

So for water it is: 418 and for copper it is 38.5

OpenStudy (aaronq):

what are those values of?

OpenStudy (anonymous):

of the heat released from the water and copper, sorry im lost on this question

OpenStudy (aaronq):

so for copper... q=(100g)(0.385 J/g–°C)(298 K-1358K) this was all given in the question

OpenStudy (anonymous):

so for copper it is: -40810?

OpenStudy (aaronq):

don't forget the units 40810 Joules

OpenStudy (aaronq):

-

OpenStudy (aaronq):

it's negative because energy was released

OpenStudy (anonymous):

ok so that is the heat released from the copper block, how do you solve the other parts of the question?

OpenStudy (aaronq):

look at this example: Heat required to convert 100 °C water to 100 °C steam q = m·ΔHv where q = heat energy m = mass ΔHv = heat of vaporization q = (25 g)x(2257 J/g) q = 56425 J Heat required to convert 100 °C water to 100 °C steam = 56425 J

OpenStudy (anonymous):

I think i may have got it could you tell me if I am right?

OpenStudy (anonymous):

4.2 grams are steam at 123 degrees Celsius, 95.8 grams are liquid at 100 degrees Celsius?

OpenStudy (aaronq):

do you have equations to go with those answers

OpenStudy (anonymous):

I used equations to get to those answers, yes.

OpenStudy (aaronq):

it looks like more water should be in the gas phase

OpenStudy (anonymous):

the calculations show that only 4.2 grams are steam, what answer did you get?

OpenStudy (anonymous):

do you know what the answer to this question is, it is really confusing: 3. Given the following equations, calculate the ΔH of reaction in which sulfur dioxide (SO2) is reduced by hydrogen disulfide (H2S) to give elemental sulfur and water (H2O).

OpenStudy (anonymous):

Does this make any sense to you?

OpenStudy (aaronq):

SO2 + H2S -> S(s) + H2O balance it then you have to do it form their enthalpies of formation ΔH = (respective coefficients)(products)-(respective coefficients)(reactants)

OpenStudy (anonymous):

i created an overall equation which is 2H2s+so2 Delta H 3S+2H2O How would you use these three equations: 1.) 2H2(g)+2s(s)2H2S(g)=-41.2 kJ 2.) 2SO2 (g)2S (s)+2O2(g)=593.6 kJ 3.) 2H2+2O2(g)2H2O(g) =-571.6 kJ

OpenStudy (anonymous):

sorry the overall equation is: 2H2s+so23S+2H2O with no delta H

OpenStudy (anonymous):

the square boxes represent the yield

OpenStudy (aaronq):

thats good, you have to reverse some to make it look like the one you need, then just add the individual enthalpies

OpenStudy (anonymous):

yeah that is where i am having trouble: would it be changing the three given: H2 (g) + S (s) → H2S (g) ΔH = –20.6 kJ SO2 (g) → S (s) + O2 (g) ΔH = +296.8 kJ H2 (g) + O2 (g) → H2O (g) ΔH = –285.8 kJ to the three i tried to fo to match the target one: 1.) 2H2(g)+2s(s)2H2S(g)=-41.2 kJ 2.) 2SO2 (g)2S (s)+2O2(g)=593.6 kJ 3.) 2H2+2O2(g)2H2O(g) =-571.6 kJ

OpenStudy (anonymous):

Not sure if i did it right

OpenStudy (aaronq):

2H2s+so23S+2H2O H2 (g) + S (s) → H2S (g) ΔH = –20.6 kJ (x2 and reverse) SO2 (g) → S (s) + O2 (g) ΔH = +296.8 kJ (x2) H2 (g) + O2 (g) → H2O (g) ΔH = –285.8 kJ (x2) does that work out?

OpenStudy (anonymous):

what do you mean when you say to reverse the first one?

OpenStudy (anonymous):

for the second to that is what i did to get: 2.) 2SO2 (g)2S (s)+2O2(g)=593.6 kJ 3.) 2H2+2O2(g)2H2O(g) =-571.6 kJ

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

not sure about the first one

OpenStudy (aaronq):

reverse it H2S (g) →H2 (g) + S (s) ΔH = + 20.6 kJ

OpenStudy (anonymous):

so when you add the entholpies you get: -1144.6?

OpenStudy (anonymous):

is that correct?

OpenStudy (aaronq):

i haven't done them myself, sorry. if you balanced the equations, that should be the right answer

OpenStudy (anonymous):

Thanks